OpenStudy (anonymous):

multiply and simplify 1/2 x [2 1/6 +7/8 really need the answer asap fan and medal

OpenStudy (anonymous):

do you mean \[\frac{1}{2} \times (2\frac{1}{6}+\frac{7}{8})\] or \[\frac{1}{2} \times (\frac{21}{6}+\frac{7}{8})\]

OpenStudy (anonymous):

the 1st one

OpenStudy (anonymous):

ok simply distribute to each term: But first make 2 and 1/6 a fraction so:\[2+1/6 = 13/6\]so the problem becomes \[\frac{1}{2} \times (\frac{13}{6}+\frac{7}{8})\] distribute the 1/2 so: \[\frac{1}{2} \times \frac{13}{6}+\frac{1}{2} \times\frac{7}{8}\]can you take it from here?

OpenStudy (anonymous):

dude i just need the answer to the question i only have 3 minutes and 4 questions left

OpenStudy (anonymous):

hmmm why not just put that problem in you calculator then?

OpenStudy (anonymous):

*your

OpenStudy (anonymous):

ok